The businesses information for 2 products is:

Product % of total sales * Break-even point.Product A

ProductB

Sales Price

$16

$29

Variable Costs

$6

$9

Annual sales (units)

12,000

8,000

Fixed Costs(all products)

$120,000

Complete the following:

1.Identify the sales mix.

2.Calculate the unit contribution margin for each product.

3.Calculate the weighted average unit contribution margin (WACM):

4.Calculate the break-even point based on the WACM

5.Calculate the break-even point for the individual products as:
Answer #1

1. sales mix = (12,000 / 8000) =>3:2.

(3 units of A and 2 units of B).

2.unit contribution margin = sale price per unit - variable cost per unit.

product A = $16-6=>$10.

product B =$29-9

=>$20.

3.weighted average contribution margin

[($10 * 3 units weight) + ($20 *2 units weight)] / (3+2)

=>(30+40) /5

=>$14.

4.break even point = 120,000 fixed costs / $14 contribution

=>8,572 units.

5.break even point for individual products

product A = 8,572 units * 3 /5 =>5,143 units.

product B = 8,572*2/5 =.3,429 units.